Hello, Assume that a company produce 100 units of laptops in assembly line process, using the...

Hello,

Assume that a company produce 100 units of laptops in assembly line process, using the bellow information, calculate the throughput ratio to measure the process efficiency:

Process time = 11 hours    ,    Waiting time = 2 hours

Could you please answer with an explaination ? As I know here we are supposed to use the throughput ratio expressed as a percentage.

Thanks in advance :)
Answer #1

Throughput ratio can be defined as the return per factory hour divided by the cost per factory hour.

which is 11/11+2 = 11/13 = 0.8461

which is equal to the process efficiency = 84.61%

The ratio means that company is paying its resources for 13 hours ( 11 productive and 2 non productive hours), but it is getting a work worth 11 hours only. Hence the efficiency of 84.61%.